# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ II-D Hybrid Systems
is said to be ISS with respect to a compact set $\mathcal{A}$ if there exist $\beta\in\mathcal{KL}$ [^1] A continuous function $\beta\colon\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ belongs to class $\mathcal{KL}$ if 1) $\beta(r,s)$ is non-decreasing in $r$, 2) $\lim_{r\to 0^{+}}\beta(r,s)=0$ for each fixed $s\in\mathbb{R}_{\geq 0}$, 3) $\beta(r,s)$ is non-increasing in $s$, and 4) $\lim_{s\to\infty}\beta(r,s)=0$ for each fixed $r\in\mathbb{R}_{\geq 0}$. and $\gamma\in\mathcal{K}$ [^2] A continuous function $\gamma\colon[0,a]\to\mathbb{R}_{\geq 0}$ with $a>0$ belongs to class $\mathcal{K}$ if 1) $\gamma$ is strictly increasing and 2) $\gamma(0)=0$. such that, for each $\phi(0,0)\in\mathbb{R}^{n}$, every solution pair $(\phi,u)$ satisfies
$|\phi(t,j)|_{\mathcal{A}}\leq\max\{\beta(|\phi(0,0)|_{\mathcal{A}},t+j),\gamma(|u|_{\infty})\}$
for all $(t,j)\in\text{dom }\phi$. $\triangle$
@@
Observe $(\phi,u)$ is a solution pair of (4) if, for each $j\in\mathbb{Z}_{\geq 0}$, the map $t\mapsto\phi(t,j)$ is locally absolutely continuous, $t\mapsto u(t,j)$ is Lebesgue measurable and locally essentially bounded over $\{t\colon(t,j)\in\text{dom}(\phi,u)\}$, dom $\text{dom }\phi=\text{dom }u$ where $\text{dom}(\phi,u)\triangleq\text{dom }\phi$ represents the hybrid time domain of $(\phi,u)$, $(\phi(t,j-1),u(t,j-1))\in C$ and $\dot{\phi}(t,j-1)=f(\phi(t,j-1),u(t,j-1))$ for almost all $t\in I_{j}$, and $(\phi(t_{j-1},j-1),u(t_{j-1},j-1))\in D$ and $\phi(t_{j-1},j)\in G(\phi(t_{j-1},j-1),u(t_{j-1},j-1))$. Note that the $\mathcal{L}_{\infty}$ norm, i.e., $|\cdot|_{\infty}$, is defined in [35, Section II].
III Problem Formulation
@@
Consider a MAS of $N\in\mathbb{Z}_{\geq 2}$ agents, which are enumerated by the elements of $\mathcal{V}$. The motion model of agent $p\in\mathcal{V}$ is
$\dot{x}_{p}=Ax_{p}+Bu_{p},\quad y_{p}=Hx_{p},$ (5)
@@
where $x_{p}\in\mathbb{R}^{n}$, $u_{p}\in\mathbb{R}^{d}$, and $y_{p}\in\mathbb{R}^{m}$ are the state variable, control input, and output of the agent, respectively. In addition, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times d}$, and $H\in\mathbb{R}^{m\times n}$ represent the known state, control effectiveness, and output matrices, respectively. To facilitate cooperation, the agents may exchange information via an intermittently available communication network, which is modeled by a static, undirected, and connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with cluster set $\mathcal{C}(\mathcal{V})$ and inter-cluster set $\mathcal{I}(\mathcal{V})$. Thus, $\mathcal{G}$ possesses $M$ cluster sub-graphs, i.e., $\{\mathcal{G}[\mathcal{V}_{p}]\}_{p\in[M]}$, and $M^{\ast}$ inter-cluster sub-graphs, i.e., $\{\mathcal{G}[\mathcal{V}^{r}]\}_{r\in[M^{\ast}]}$. Further, for each $p\in\mathcal{V}$, agent $p$ can measure the outputs of its neighbors and itself, that is, agent $p$ can measure $\{y_{q}\}_{q\in\mathcal{N}_{p}\cup\{p\}}$.
@@
The objective of this work is to develop a distributed controller for each agent $p\in\mathcal{V}$ that yields consensus in the agents’ states, that is, $x_{p}=x_{q}$ for all distinct $p,q\in\mathcal{V}$. In addition, for each $p\in\mathcal{V}$, the controller of agent $p$ should only utilize intermittent measurements of the outputs in $\{y_{q}\}_{q\in\mathcal{N}_{p}\cup\{p\}}$, which may be obtained asynchronously. To be more precise, the measurements of outputs belonging to agents in $\mathcal{N}_{p}\cap\mathcal{C}_{p}$, $\mathcal{N}_{p}\cap\mathcal{V}^{r}$, and $\mathcal{N}_{p}\cap\mathcal{V}^{s}$, for $r\neq s\in[M^{\ast}]$ with $\mathcal{N}_{p}\cap\mathcal{V}^{r}\neq\varnothing$ and $\mathcal{N}_{p}\cap\mathcal{V}^{s}\neq\varnothing$, may occur at different moments in time and at different rates. Such a control strategy better leverages limited resources relative to continuous communication and/or sensing alternatives, readily integrates in digital hardware, and accommodates the natural asynchronous flow of information within MASs.
IV Hybrid System Development
For each $p\in\mathcal{V}$, let $0<T_{1}^{p}\leq T_{2}^{p}$ be user-defined constants. Furthermore, let $\tau_{p}\in[0,T_{2}^{p}]$ be a timer for agent $p$ that evolves according to the hybrid system
@@ IV Hybrid System Development
with initial condition $\tau_{p}(0,0)\in[0,T_{2}^{p}]$. [^3] Examples that leverage the timer mechanism in (6) can be found in [17], [32], and [36]. The hybrid dynamics in (6) enables the construction of increasing sequences of time, e.g., $\{t_{k}^{p}\}_{k=0}^{\infty}$, given a complete solution $\phi_{\tau_{p}}$, [^4] Note, $\phi_{\tau_{p}}$ represents the value of $\tau_{p}$ along a solution $\phi$ of (6). where the event time $t_{k}^{p}$ represents the $k^{\text{th}}$ instant $\tau_{p}=0$. Moreover, the event times satisfy the following inequalities: for all $k\in\mathbb{Z}_{\geq 1}$,
$T_{1}^{p}\leq t_{k+1}^{p}-t_{k}^{p}\leq T_{2}^{p}.$
Likewise, for every inter-cluster $r\in[M^{\ast}]$, let $0<T_{3}^{r}\leq T_{4}^{r}$ be user-defined constants, and let $\rho_{r}\in[0,T_{4}^{r}]$ be a timer for $\mathcal{V}^{r}$ that evolves according to the hybrid system
@@
with initial condition $\rho_{r}(0,0)\in[0,T_{4}^{r}]$. The hybrid dynamics in (7) enables the construction of increasing sequences of time, e.g., $\{\mathtt{t}_{k}^{r}\}_{k=0}^{\infty}$, given a complete solution $\phi_{\rho_{r}}$, where the event time $\mathtt{t}_{k}^{r}$ represents the $k^{\text{th}}$ instant $\rho_{r}=0$. Moreover, the event times satisfy the following inequalities: for all $k\in\mathbb{Z}_{\geq 1}$,
$T_{3}^{r}\leq\mathtt{t}_{k+1}^{r}-\mathtt{t}_{k}^{r}\leq T_{4}^{r}.$
These timers can be treated as independent autonomous systems that trigger certain actions. In particular, $\tau_{p}$ will be used to dictate when agent $p$ obtains output information from same-cluster neighbors, and $\rho_{r}$ will be used to determine when the agents in inter-cluster $r$ obtain output information from same-inter-cluster neighbors.
For every agent $p\in\mathcal{V}$ and inter-cluster $r\in[M^{\ast}]$, let $\eta_{p}\in\mathbb{R}^{m}$ and $\zeta_{pr}\in\mathbb{R}^{m}$ be auxiliary variables that evolve according to
@@
respectively, where $K_{\eta,p},K_{\zeta,pr}\in\mathbb{R}^{m\times m}$ are user-defined matrices. Loosely speaking, a solution to (6) and (8) would yield an approximation of the consensus-like term $\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p})$ given continuous-time trajectories for $\{y_{q}\}_{q\in(\mathcal{N}_{p}\cap\mathcal{C}_{p})\cup\{p\}}$. The approximation is intermittently reset to the corresponding value of $\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p})$ and allowed to flow in an open-loop fashion according to the flow equation in (8). Note that if $0<T_{1}^{p}<T_{2}^{p}$, then the times when $\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p})$ is sampled are intermittent since the timer may be reset to any number in the interval $[T_{1}^{p},T_{2}^{p}]$. Periodic sampling can be created by selecting $0<T_{1}^{p}=T_{2}^{p}$. An analogous comment follows for (7) and (9).
@@
Given a user-defined matrix $K_{u}\in\mathbb{R}^{d\times m}$, the controller of agent $p$ is
$u_{p}\triangleq K_{u}\Bigg{(}\eta_{p}+\sum_{r\in[M^{\ast}]}\zeta_{pr}\Bigg{)}\in\mathbb{R}^{d}.$ (10)
@@
By construction, the controller in (10) is distributed as it only requires output information from neighboring agents and the implementing agent itself. Also, the output measurements are obtained intermittently, provided $T_{1}^{p}<T_{2}^{p}$ and $T_{3}^{r}<T_{4}^{r}$ for all $r\in[M^{\ast}]$, and the output measurements belonging to agents in $\mathcal{N}_{p}\cap\mathcal{C}_{p}$, $\mathcal{N}_{p}\cap\mathcal{V}^{r}$, and $\mathcal{N}_{p}\cap\mathcal{V}^{s}$, for $r\neq s\in[M^{\ast}]$, may be obtained asynchronously. Note the timer $\tau_{p}$ determines when agent $p$ samples the data in $\{y_{q}-y_{p}\}_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}$ and updates the $\eta_{p}$ component of its controller given (8). Likewise, the timer $\rho_{r}$ determines when agent $p$ samples the data in $\{y_{q}-y_{p}\}_{q\in\mathcal{N}_{p}\cap\mathcal{V}^{r}}$ and updates the $\zeta_{pr}$ component of its controller given (9). If agent $p$ is not a member of inter-cluster $r$ (i.e., $p\notin\mathcal{V}^{r}$), $\zeta_{pr}\equiv 0_{m}$. To facilitate the derivation of the closed-loop ensemble hybrid system, let
$\tilde{\eta}_{p}$ $\triangleq\eta_{p}-\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p})\in\mathbb{R}^{m},$ (11)
$\tilde{\zeta}_{pr}$ $\triangleq\zeta_{pr}-\sum_{q\in\mathcal{N}_{p}\cap\mathcal{V}^{r}}(y_{q}-y_{p})\in\mathbb{R}^{m}.$ (12)
We provide the following notation to aid the writing of concise expressions. Let
$x$ $\triangleq(x_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{nN},$ $\eta$ $\triangleq(\eta_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$
$\zeta_{r}$ $\triangleq(\zeta_{pr})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$ $\zeta$ $\triangleq(\zeta_{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{mNM^{\ast}},$
$\tilde{\eta}$ $\triangleq(\tilde{\eta}_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$ $\tilde{\zeta}_{r}$ $\triangleq(\tilde{\zeta}_{pr})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$
$\tilde{\zeta}$ $\triangleq(\tilde{\zeta}_{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{mNM^{\ast}},$ $\tau$ $\triangleq(\tau_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{N},$
$\rho$ $\triangleq(\rho_{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{M^{\ast}}.$

### v3

@@ 2.4 Hybrid Systems
is said to be input-to-state stable (ISS) with respect to a compact set $\mathcal{A}$ if there exist $\beta\in\mathcal{KL}$ [^4] A continuous function $\beta\colon\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ belongs to class $\mathcal{KL}$ if 1) $\beta(r,s)$ is non-decreasing in $r$, 2) $\lim_{r\to 0^{+}}\beta(r,s)=0$ for each fixed $s\in\mathbb{R}_{\geq 0}$, 3) $\beta(r,s)$ is non-increasing in $s$, and 4) $\lim_{s\to\infty}\beta(r,s)=0$ for each fixed $r\in\mathbb{R}_{\geq 0}$. and $\gamma\in\mathcal{K}$ [^5] A continuous function $\gamma\colon[0,a]\to\mathbb{R}_{\geq 0}$ with $a>0$ belongs to class $\mathcal{K}$ if 1) $\gamma$ is strictly increasing and 2) $\gamma(0)=0$. such that, for each $\phi(0,0)\in\mathbb{R}^{n}$, every solution pair $(\phi,{d})$ satisfies
$|\phi(t,j)|_{\mathcal{A}}\leq\max\{\beta(|\phi(0,0)|_{\mathcal{A}},t+j),\gamma(|{d}|_{\infty})\}$
for all $(t,j)\in{\text{dom }(\phi,d)}$. $\triangle$
@@
Note, $(\phi,{d})$ is said to be a solution pair of (4) if, for each $j\in\mathbb{Z}_{\geq 0}$, the mapping $t\mapsto\phi(t,j)$ is locally absolutely continuous, $t\mapsto{d}(t,j)$ is Lebesgue measurable and locally essentially bounded over $\{t\colon(t,j)\in\text{dom}(\phi,{d})\}$, $\text{dom }\phi=\text{dom }{d}$ where $\text{dom}(\phi,{d})\coloneqq\text{dom }\phi$ represents the hybrid time domain of $(\phi,{d})$, $(\phi(t,j-1),{d}(t,j-1))\in C$ and $\dot{\phi}(t,j-1)=f(\phi(t,j-1),{d}(t,j-1))$ for almost all $t\in I_{j}$, and $(\phi(t_{j-1},j-1),{d}(t_{j-1},j-1))\in D$ and $\phi(t_{j-1},j)\in G(\phi(t_{j-1},j-1),{d}(t_{j-1},j-1))$. Note, $|\cdot|_{\infty}$ is defined in [Cai.Teel2009, Right Column of Page 48]. The class of perturbations defined within Definition 2 will serve as the non-vanishing perturbations discussed in Section 6.
3 Problem Formulation
@@
Consider a MAS of $N\in\mathbb{Z}_{\geq 2}$ agents indexed by the elements of $\mathcal{V}$, where the dynamics of agent $p\in\mathcal{V}$ is
$\dot{x}_{p}=Ax_{p}+Bu_{p},\quad y_{p}=Hx_{p}.$ (5)
@@
Note, $x_{p}\in\mathbb{R}^{n}$, $u_{p}\in\mathbb{R}^{d}$, and $y_{p}\in\mathbb{R}^{m}$ represent the state variable, control input, and output of the agent, respectively. Furthermore, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times d}$, and $H\in\mathbb{R}^{m\times n}$ represent the known state, control effectiveness, and output matrices, respectively. To facilitate cooperation, the agents may exchange information via a communication network modeled by a static, undirected, and connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Depending on the application, a MAS may need to be divided into several smaller teams, which may require different communication rates. As a result, the node set $\mathcal{V}$ of $\mathcal{G}$ is partitioned to yield $M$ cluster sub-graphs, i.e., $\{\mathcal{G}[\mathcal{V}_{p}]\}_{p\in[M]}$, and $M^{\ast}$ inter-cluster sub-graphs, i.e., $\{\mathcal{G}[\mathcal{V}^{r}]\}_{r\in[M^{\ast}]}$. [^6] Please see Section 2.3 for the precise details regarding the cluster and inter-cluster sub-graphs considered in this work. For each $p\in\mathcal{V}$, agent $p$ can measure the outputs of its same-cluster neighbors and itself, that is, $\{y_{q}\}_{q\in(\mathcal{N}_{p}\cap\mathcal{C}_{p})\cup\{p\}}$, at isolated times defined by the sequence $\{t_{k}^{p}\}_{k=0}^{\infty}$ which satisfies
$T_{1}^{p}\leq t_{k+1}^{p}-t_{k}^{p}\leq T_{2}^{p}$ (6)
for all $k\in\mathbb{Z}_{\geq 1}$. Note, the parameters $T_{1}^{p}$ and $T_{2}^{p}$ should be selected to satisfy $0<T_{1}^{p}\leq T_{2}^{p}$. Recall, the assigned cluster of agent $p$ is denoted by $\mathcal{C}_{p}$ (see Section 2.3). If agent $p$ is also a member of inter-cluster $r\in[M^{\ast}]$, then agent $p$ can also measure the output of its same-$r$-inter-cluster neighbors and itself at isolated times defined by the sequence $\{\mathtt{t}_{k}^{r}\}_{k=0}^{\infty}$ which satisfies
$T_{3}^{r}\leq\mathtt{t}_{k+1}^{r}-\mathtt{t}_{k}^{r}\leq T_{4}^{r}$ (7)
for all $k\in\mathbb{Z}_{\geq 1}$. Note, the parameters $T_{3}^{r}$ and $T_{4}^{r}$ should be selected to satisfy $0<T_{3}^{r}\leq T_{4}^{r}$. In light of (6) and (7), the rates of output measurements in the cluster of agent $p$ and inter-cluster $r$ can be controlled through the selection of the parameters $T_{1}^{p}$, $T_{2}^{p}$, $T_{3}^{r}$, and $T_{4}^{r}$.
The objective of this work is to develop a distributed controller for each agent $p\in\mathcal{V}$ that yields consensus in the agents’ states, that is,
$x_{p}=x_{q}$
@@
for all $p,q\in\mathcal{V}$ with stable behavior and robustness. In addition, for each $p\in\mathcal{V}$, the controller of agent $p$ should only utilize intermittent measurements of the outputs in $\{y_{q}\}_{q\in\mathcal{N}_{p}\cup\{p\}}$, which may be obtained asynchronously. Such a control strategy better leverages limited resources relative to continuous communication and/or sensing alternatives, readily integrates in digital hardware, and accommodates the natural asynchronous flow of information within MASs.
4 Hybrid System Modeling
Within this section, we develop the closed-loop ensemble hybrid system under our proposed distributed consensus controller. We begin by introducing the timer mechanisms responsible for the generation of the sequences $\{t_{k}^{p}\}_{k=0}^{\infty}$ and $\{\mathtt{t}_{k}^{r}\}_{k=0}^{\infty}$ discussed in the previous section. Two different hybrid estimators are introduced next, which are the foundation of our distributed consensus controller. The final sub-section presents auxiliary error systems and ensemble variables that support the derivation of the closed-loop ensemble hybrid system.
4.1 Timer Mechanisms
Let $0<T_{1}^{p}\leq T_{2}^{p}$ be user-defined constants for each $p\in\mathcal{V}$, where $T_{1}^{p}$ can be used to account for the sampling rate of the slowest sensor of agent $p$ and $T_{2}^{p}$ can be used to account for the MATI of agent $p$. Furthermore, let $\tau_{p}\in[0,T_{2}^{p}]$ denote the time kept by a timer mechanism belonging to agent $p$ whose evolution is modeled by the hybrid system
@@ 4.1 Timer Mechanisms
with initial condition $\tau_{p}(0,0)\in[0,T_{2}^{p}]$. [^7] Examples that leverage the timer mechanism in (8) can be found in [Phillips.Sanfelice2019a], [Goebel.Sanfelice.ea2012], and [Zegers.Phillips.ea2024]. The hybrid dynamics in (8) and the onboard clock of agent $p$ enable the construction of increasing sequences of event times by using the following procedure: 1) set $t_{k}^{p}$ for $k=0$ to the initial time given by the onboard clock of agent $p$, 2) increment $k$ by one and allow the timer of agent $p$ to flow using the flow equation in (8) and the initial condition $\tau_{p}(0,0)\in[0,T_{2}^{p}]$ (during this step, the value of $\tau_{p}$ is made to decrease at a rate of 1-1 second/second from its initial value), 3) once $\tau_{p}=0$, set $t_{k}^{p}$ for $k=1$ to the current time kept by the onboard clock of agent $p$ and reset $\tau_{p}$ to any element in $T_{1}^{p},T_{2}^{p}]$ (that is, allow the timer of agent $p$ to jump according to the jump inclusion in (8)), 4) increment $k$ by one and allow the timer of agent $p$ to flow using the flow equation in (8) and the value of $\tau_{p}$ after the jump from Step 3), 5) repeat Steps 3) and 4) in aeternum. The result of this procedure is an increasing sequence of event times, where the event time $t_{k}^{p}$ represents the $k^{\text{th}}$ instant $\tau_{p}=0$ for agent $p$. Note, assuming the onboard clock of agent $p$ increases at a constant rate of 1 second/second, the sequence of event times generated by (8) and the aforementioned procedure satisfies (6) for all $k\in\mathbb{Z}_{\geq 1}$.
Likewise, for every inter-cluster $r\in[M^{\ast}]$, let $0<T_{3}^{r}\leq T_{4}^{r}$ be user-defined constants playing a similar role to those of $T_{1}^{p}$ and $T_{2}^{p}$, and let $\rho_{r}\in[0,T_{4}^{r}]$ denote the time kept by a timer mechanism assigned to inter-cluster $r$ that evolves according to the hybrid system
@@
with initial condition $\rho_{r}(0,0)\in[0,T_{4}^{r}]$. The hybrid dynamics in (9) enables the construction of increasing sequences of event times, e.g., $\{\mathtt{t}_{k}^{r}\}_{k=0}^{\infty}$ following a similar procedure to the one described above, where the event time $\mathtt{t}_{k}^{r}$ represents the $k^{\text{th}}$ instant $\rho_{r}=0$. Note, the event times in $\{\mathtt{t}_{k}^{r}\}_{k=0}^{\infty}$ can be defined by using a common clock available to all agents of inter-cluster $r$. For simplicity, we assume the onboard clocks of all agents are synchronized, which facilitates the construction of event time sequences for agents and inter-clusters defined relative to a common time standard. An exemplar result enabling distributed clock synchronization is [Guarro.Sanfelice2023]. Note, assuming the onboard clocks of the agents are synchronized and increase at a constant rate of 1 second/second, the sequence of event times generated by (9) and a sequence building procedure similar to the one discussed above satisfies (7) for all $k\in\mathbb{Z}_{\geq 1}$. A convenient feature of the hybrid system in (8) is that intermittent event times can be generated by ensuring $0<T_{1}^{p}<T_{2}^{p}$ holds. Nonetheless, (8) can produce periodic event times provided $0<T_{1}^{p}=T_{2}^{p}$ holds. An analogous comment follows for the event times generated by the hybrid system in (9). The timers modeled by (8) and (9) can be treated as independent autonomous systems capable of triggering intermittent and asynchronous actions as defined next.
4.2 Consensus Controller Design
To achieve consensus, we design a first-order hybrid controller that utilizes output measurements from connected agents, when they are available, to drive the state of each agent with dynamics in (5). For every agent $p\in\mathcal{V}$ and inter-cluster $r\in[M^{\ast}]$, let $\eta_{p}\in\mathbb{R}^{m}$ and $\zeta_{pr}\in\mathbb{R}^{m}$ be auxiliary variables that evolve according to
@@ 4.2 Consensus Controller Design
respectively, where $K_{\eta,p},K_{\zeta,pr}\in\mathbb{R}^{m\times m}$ are user-defined matrices. The systems in (10) and (11) model estimators belonging to agent $p$ that are intermittently reset with output information during certain jumps and allowed to evolve in an open-loop fashion during flows. In particular, when a jump is caused by $\tau_{p}=0$, the value of $\eta_{p}$ is reset with the value of a consensus-like term computed using only output information from the same-cluster neighbors of agent $p$, as defined by the jump equation in (10). When $\tau_{p}\in[0,T_{2}^{p}]$, the flow equation (i.e., LTI system) in (10) defines an open-loop trajectory, that is, a trajectory not influenced by output feedback. Similarly, when a jump is caused by $\rho_{r}=0$, the value of $\zeta_{pr}$ is reset with the value of a consensus-like term computed using only output information from the neighbors of agent $p$ in inter-cluster $r$, as defined by the jump equation in (11). When $\rho_{r}\in[0,T_{4}^{r}]$, the flow equation (i.e., LTI system) in (11) defines an open-loop trajectory. Since the timers modeled by (8) and (9) are independent, the resets experienced by $\eta_{p}$ and $\zeta_{pr}$ may be asynchronous. It is worth noting that the hybrid dynamics in (10) and (11) for every $p\in\mathcal{V}$ and $r\in[M^{\ast}]$ induce information flows between agents that are intermittent, asynchronous, and directed, which are taking place over a static, connected, and undirected graph. Moreover, at consensus, i.e., when $x_{p}=x_{q}$ for all $p,q\in\mathcal{V}$, the variables $\eta_{p}$ and $\zeta_{pr}$ both equal $0_{n}$ given (5), (10), and (11). It is also worth noting that, even though the estimators values $\eta_{p}$ and $\zeta_{pr}$ evolve in an open-loop fashion during flows, the timer parameters $T_{2}^{p}$ and $T_{4}^{r}$ bound the period of respective flow and will be selected to ensure the stability of the closed-loop system.
@@
Given a user-defined matrix $K_{u}\in\mathbb{R}^{d\times m}$ and the estimates produced by the estimators in (10) and (11), the controller of agent $p$ is designed as
$u_{p}\coloneqq K_{u}\Bigg(\eta_{p}+\sum_{r\in[M^{\ast}]}\zeta_{pr}\Bigg)\in\mathbb{R}^{d}.$ (12)
@@
By construction, the controller in (12) is distributed as it only requires output information from neighboring agents and the implementing agent itself. The output measurements can be obtained intermittently provided $T_{1}^{p}<T_{2}^{p}$ and $T_{3}^{r}<T_{4}^{r}$ for all inter-clusters $r\in[M^{\ast}]$. Furthermore, the output measurements belonging to agents in $\mathcal{N}_{p}\cap\mathcal{C}_{p}$, $\mathcal{N}_{p}\cap\mathcal{V}^{r}$, and $\mathcal{N}_{p}\cap\mathcal{V}^{s}$, where $r\neq s\in[M^{\ast}]$, may be obtained asynchronously given the independence of the timer mechanisms. Note, if agent $p$ is not a member of inter-cluster $r$ (i.e., $p\notin\mathcal{V}^{r}$), $\zeta_{pr}\equiv 0_{m}$.
4.3 Hybrid System Modeling
To facilitate the derivation of the closed-loop ensemble hybrid system, we will leverage the errors systems
$\tilde{\eta}_{p}$ $\coloneqq\eta_{p}-\sum_{q\in\mathcal{N}_{p}\cap\mathcal{C}_{p}}(y_{q}-y_{p})\in\mathbb{R}^{m},$ (13)
$\tilde{\zeta}_{pr}$ $\coloneqq\zeta_{pr}-\sum_{q\in\mathcal{N}_{p}\cap\mathcal{V}^{r}}(y_{q}-y_{p})\in\mathbb{R}^{m}.$ (14)
Furthermore, we provide the following notation to aid the writing of concise expressions. Let
$x$ $\coloneqq(x_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{nN},$ $\eta$ $\coloneqq(\eta_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$ $\zeta_{r}$ $\coloneqq(\zeta_{pr})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$ $\zeta$ $\coloneqq(\zeta_{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{mNM^{\ast}},$
$\tilde{\eta}$ $\coloneqq(\tilde{\eta}_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$ $\tilde{\zeta}_{r}$ $\coloneqq(\tilde{\zeta}_{pr})_{p\in\mathcal{V}}\in\mathbb{R}^{mN},$ $\tilde{\zeta}$ $\coloneqq(\tilde{\zeta}_{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{mNM^{\ast}},$ $\tau$ $\coloneqq(\tau_{p})_{p\in\mathcal{V}}\in\mathbb{R}^{N},$
$\rho$ $\coloneqq(\rho_{r})_{r\in[M^{\ast}]}\in\mathbb{R}^{M^{\ast}}.$